Tarek has been thinking about getting a credit card, but he is skeptical that a credit card can be a useful spending tool for him. Instead, Tarek has always used a debit card. His reasoning is that using a debit card provides protection from overspending; after all, he can only spend what is in his account.
Assume Tarek now has a credit card and he carries a $2,000monthly balance. If the credit card company uses a 4% multiplier to determine the minimum monthly payment and the card has a 15% APR and 30-day billing cycle, what will be the minimum monthly payment? How much of the first payment will go towards paying interest?

The minimum monthly payment will be $ .

Amount of first payment that will go towards paying interest is $ .
Answer #1

Monthly balance = $2,000

Using 4% multiplier

Annual percentage rate = 15% = 15/12 = 1.25%

First payment towards paying interest = $2,000*1.25% = $24

Hence, tarek has to owe = $2,000 + $25 = $2,025

Minimum monthly payment = $2025 * 4% = $81